Case Study - Laboratory
- Industry: Research & Development
- Type: Diagnostic
- Square Footage: 350,000
- Location: Undisclosed
- View Comprehensive Report
This report pertains to a heat recovery unit at an R&D laboratory. This unit was reported as not operating in a stable condition. It was further explained that the BMS cannot provide a stable negative static pressure condition in the exhaust duct upstream of the HEPA filter bank. This unit is equipped with a VI as the volume control method. This report outlines the observations and steps taken to attempt to correct the problem.
The VI was reported on this unit to be hunting. Examined the VI, controller, linkage, etc., and determined that the VI was hunting at approximately 3.7 VDC at the electric to pneumatic transducer (EP). Measured the command voltage at the controller and at the EP and found that the voltage at the EP was hunting while the controller output was not. Using the same wire, we connected a signal generator and found the same condition. Installed a temporary cable outside of the conduit and found a stable signal at both ends with the signal generator and the BAS contractor controller. Traced the problem to a ground loop which was causing interference in the conduit Disconnected the cable shield drain wire and the problem ceased. The cable should be replaced.
The best solution would be to retrofit each system with a VFD and allow the BAS contractor to control fan speed to satisfy the flow and pressure requirements. If VFDs are not an option, then each fan system should be tested to determine the point of instability and then the fans should be re-sheaved to a speed below the instability range. Additional testing and/or sheave changes may be necessary as the systems are balance.
Please see our report data, conclusions and recommendations.